Given the equation 5 + x - 12 = x - 7.
Part A. Solve the equation 5 + x - 12 = x - 7. In your final answer, be sure to state the solution and include all of your work.
Part B. Use the values x = -4, 0, 5 to prove your solution to the equation 5 + x - 12 = x - 7. In your final answer, include all of your calculations.

Answer by richard1234(7193)

A. Subtract x from both sides: (i.e. 5+x-12 = x-7 <--> 5-12 = -7)

This equation is identically true, so it holds no matter what x is.

B. This one is pretty self-explanatory.